OpenStudy (anonymous):

What is the axis of symmetry for the function? y = x2 + 3x - 4 Answers: a) y = -2 b) y = -1.5 c) x = -2 d) x = -1.5

jimthompson5910 (jim_thompson5910):

hint: Axis of Symmetry for y = ax^2 + bx + c is x = -b/(2a)

OpenStudy (anonymous):

O.o i don't get it D:

jimthompson5910 (jim_thompson5910):

what are a, b and c in this case?

OpenStudy (anonymous):

x2, 3x, -4

OpenStudy (anonymous):

i think...

jimthompson5910 (jim_thompson5910):

a, b, and c are the numbers without the variables

jimthompson5910 (jim_thompson5910):

so, a = 1 b = 3 c = -4

OpenStudy (anonymous):

oh, okay

jimthompson5910 (jim_thompson5910):

so what is x = -b/(2a)

OpenStudy (anonymous):

-3/(2*1)

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

that for sure wrong

jimthompson5910 (jim_thompson5910):

nope it's correct, keep going

jimthompson5910 (jim_thompson5910):

you're doubting yourself for no reason

OpenStudy (anonymous):

I think it would equal -1

OpenStudy (anonymous):

because -3 divided by 3 = -1

jimthompson5910 (jim_thompson5910):

no, it's not -3/3 it's -3/2 = ???

OpenStudy (anonymous):

omg 2*1 im stupid lol

jimthompson5910 (jim_thompson5910):

nah you're not, silly typo

OpenStudy (anonymous):

-1.5 then

jimthompson5910 (jim_thompson5910):

yep you got it

jimthompson5910 (jim_thompson5910):

so x = -1.5

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

you're welcome